MICROWAVE SPECTROSCOPY OF GROUP-10-METAL MONOCHALCOGENIDES.

\underline{T. OKABAYASHI}, Center for Instrumental Research, Shizuoka University, Oya 836, Shizuoka 422-8529, Japan; T. KURAHARA, T. YAMAMOTO, E. Y. OKABAYASHI, and M. TANIMOTO, Department of Chemistry, Faculty of Science, Shizuoka University, Oya 836, Shizuoka 422-8529, Japan.

Group 10 metals (Ni, Pd, and Pt) and their alloys are widely used in various chemical processes as catalysts, such as the three-way catalyst and desulfurization catalyst. Since oxygen or sulfur atoms in reactants can interact with metallic atoms on the surface of such catalysts, diatomic group-10-metal monochalcogenides are often attracting interest as one of the simplest models for the reactant-catalyst interactions. Recently, the diatomic species, NiO, PtO, and PtS, have been studied by a few groups using microwave spectroscopy to obtain their detailed physico-chemical properties. In this contribution, we report on microwave spectroscopic study of two other group-10-metal monochalcogenides, PdO and NiS, generated by sputtering reaction.